OpenStudy (anonymous):

-2x + 3y = 8 3x – y = -5

OpenStudy (anonymous):

system of equations

OpenStudy (anonymous):

We could try to isolate y 3x -y = -5 3x = -5 + y 3x + 5 = y Then we substitute for y in the first. -2x + 3(3x+5) = -2x +9x + 15 = 7x + 15 7x + 15 = 8 7x = 8-15 = -7 x = -7 / 7 = -1. Now we use this in the second one. 3*-1 - y = -5 -3 -y = -5 -y = -2 y = 2

OpenStudy (anonymous):

Here is what I would do. Get the equation "3x – y = -5" and solve it for "y" Then plug in this equation into the first equation.